TREX: Logbook Entries

TREX: Site central Messages, 108 Entries..

Return to Logbook Contents Page
Entry Date Title Site Author #Graphics
265 Sun 30-Apr-2006Inverter failure ends hot-film run at 1000 LT 30 Aprcentralpoulos
262 Sun 30-Apr-2006H2 levelcentralgolubieski
258 Sat 29-Apr-2006ISFF media daycentralpoulos
256 Fri 28-Apr-2006Central Guy Tensionscentralgolubieski
255 Fri 28-Apr-2006Krypton Cleaningcentralgolubieski
243 Tue 25-Apr-2006Wireless fiddle testscentralpoulos
242 Tue 25-Apr-2006Wireless intermittent outages, return to the original configuration 12.40 centralpoulos
240 Mon 24-Apr-2006Hotfilm runcentralmilitzer
239 Mon 24-Apr-2006Cleaned 5m Kryptoncentralmilitzer
238 Mon 24-Apr-2006Rain Check, Download Datacentralmilitzer
237 Mon 24-Apr-2006Tsoil Coefficients, etc.centralmilitzer
236 Mon 24-Apr-2006Trime Measurements, Gravimetriccentralmilitzer
232 Sun 23-Apr-2006Hydrogen replaced at Centralcentralpoulos
228 Sat 22-Apr-2006Boom angles, centralcentralpoulos
225 Sat 22-Apr-2006Outages continue; New Eant at Centralcentralmilitzer
223 Fri 21-Apr-2006GPS coordinates of the hot-film sonic anemometerscentralpoulos
222 Fri 21-Apr-2006Pictures around Licorcentralpoulos
221 Fri 21-Apr-2006Central pocketec downloadcentralpoulos
219 Thu 20-Apr-2006hotfilm runningcentralmilitzer
218 Thu 20-Apr-20065m Krypton cleanedcentralmilitzer
213 Wed 19-Apr-2006Trime Measurement, Gravimetriccentralmilitzer
212 Wed 19-Apr-2006Rad.cleanedcentralmilitzer
204 Tue 18-Apr-2006hotfilm system clock aheadcentralmaclean
201 Sun 16-Apr-2006Started hotfilmcentralmaclean
197 Sat 15-Apr-2006hotfilm powercentralmaclean
196 Sat 15-Apr-2006Short local storage outagecentralmilitzer
195 Sat 15-Apr-2006Reconfigured power to hotfilm tower.centralmilitzer
194 Sat 15-Apr-200610m Rad cleaned, photoscentralpoulos
193 Sat 15-Apr-20065m TRH filter checkcentralmilitzer
191 Fri 14-Apr-2006hotfilm data system testcentralpoulos
190 Fri 14-Apr-2006Central 2m sonic, power usagecentralpoulos
186 Thu 13-Apr-2006hotfilm data system testcentralmaclean
185 Thu 13-Apr-20062m Sonic downcentralmaclean
182 Thu 13-Apr-2006Update on 35W Central Fuel Cell centralmilitzer
181 Thu 13-Apr-2006Guy Wires added to 5m towercentralmilitzer
180 Thu 13-Apr-200630m krypton cleanedcentralmilitzer
178 Wed 12-Apr-2006Update on Central Fuel Cell / Battery Statuscentralmilitzer
177 Wed 12-Apr-2006'cleaned' 10m K&Zcentralmilitzer
176 Wed 12-Apr-20065m krypton cleanedcentralmilitzer
174 Tue 11-Apr-2006Update on Central Fuel Cell / Battery Statuscentralmilitzer
171 Mon 10-Apr-2006Hotfilm runcentralmaclean
168 Sun 09-Apr-2006Update on Central Fuel Cell / Battery Statuscentralmilitzer
165 Sun 09-Apr-2006Radiometers 'Cleaned'centralmilitzer
161 Sun 09-Apr-2006HotFilm Run Sa. 4/8/06 - Su. 4/9/06centralmilitzer
160 Fri 07-Apr-2006gravimetric soil moisturecentralmilitzer
159 Fri 07-Apr-2006Central Battery Statuscentralmilitzer
158 Fri 07-Apr-2006Imitation Hotfilm Test Runningcentralmilitzer
157 Fri 07-Apr-2006Hotfilm Shutdowncentralmilitzer
155 Thu 06-Apr-2006no rain data at centralcentralhorst
154 Thu 06-Apr-2006wireless network from central downcentralhorst
152 Wed 05-Apr-2006Hotfilm runningcentralsemmer
151 Wed 05-Apr-2006central communications lostcentralgolubieski
150 Tue 04-Apr-2006ZigBee now power on/off at Centralcentralsemmer
147 Tue 04-Apr-2006lost rf connection to centralcentralhorst
145 Tue 04-Apr-2006Link to central downcentralsemmer
142 Mon 03-Apr-2006retensioned guy wirescentralhorst
141 Mon 03-Apr-2006hot film procedurescentralsemmer
136 Sun 02-Apr-2006Hot film statuscentralsemmer
134 Sat 01-Apr-200625 m sonic at centralcentralhorst
132 Sat 01-Apr-2006maintenance at centralcentralhorst
127 Thu 30-Mar-2006HotFilm shutoffcentralsemmer
124 Wed 29-Mar-2006Hot Film statuscentralsemmer
123 Wed 29-Mar-2006Logger #1 at centralcentralsemmer
121 Wed 29-Mar-2006Shutdown pyg sensor at central, logger ID 2centralsemmer
119 Tue 28-Mar-2006Hot film statuscentralsemmer
118 Tue 28-Mar-2006Hot film sonics on central nidscentralsemmer
117 Tue 28-Mar-2006hot-film sonics added to NIDScentralhorst
115 Mon 27-Mar-2006HF site back up collecting sonic data at 15:30centralsemmer
113 Mon 27-Mar-2006Battery changed at logger #1 at centralcentralsemmer
112 Mon 27-Mar-2006Hot film statuscentralsemmer
110 Sun 26-Mar-2006Deploying hot film probescentralhorst
106 Sat 25-Mar-2006hotfilm updatecentralsemmer
103 Fri 24-Mar-2006hotfilm statuscentralsemmer
101 Fri 24-Mar-2006data link to central fixedcentralhorst
99 Fri 24-Mar-2006Hotfilm work from yesterdaycentralsemmer
97 Thu 23-Mar-2006maintenance at central sitecentralhorst
94 Thu 23-Mar-2006hotfilmcentralsemmer
91 Wed 22-Mar-2006htofilm - probe position problemcentralsemmer
90 Wed 22-Mar-2006Installed hot film sensorscentralhorst
87 Tue 21-Mar-2006hotfilmcentralsemmer
86 Tue 21-Mar-2006hotfilm statuscentralsemmer
83 Mon 20-Mar-2006Htofilm statuscentralsemmer
79 Mon 20-Mar-2006hotfilmcentralsemmer
78 Sun 19-Mar-2006fuel cellcentralsemmer
77 Sun 19-Mar-2006hotfilm statuscentralsemmer
74 Sat 18-Mar-2006central fuel cellcentralsemmer
73 Sat 18-Mar-2006Central visitcentralsemmer
71 Sat 18-Mar-2006Fuel Cell at centralcentralsemmer
69 Fri 17-Mar-2006HotFilmcentralsemmer
68 Fri 17-Mar-2006Removed hot film data system and 1 m sonic centralhorst
67 Fri 17-Mar-200660 watt fuel cellcentralsemmer
66 Fri 17-Mar-2006single pyg unitscentralsemmer
62 Thu 16-Mar-2006Central visitcentraloncley
59 Thu 16-Mar-2006prometheus config changecentraloncley
57 Wed 15-Mar-2006Service visit to Central towercentralhorst
54 Wed 15-Mar-2006hotfilm analogs diedcentraloncley
53 Tue 14-Mar-2006cylinder readingscentraloncley
50 Tue 14-Mar-2006Tcase working at centralcentraloncley
47 Mon wiring errorcentraloncley
45 Sun 12-Mar-2006central died last nightcentraloncley
43 Sun 12-Mar-2006hotfilm stuff installedcentraloncley
39 Sat 11-Mar-2006visitcentraloncley
38 Sat 11-Mar-2006added fuel cellcentraloncley
36 Fri 10-Mar-2006Radiometer Power Issuescentralagee
28 Mon 06-Mar-2006central powercentraloncley
18 Wed 01-Mar-2006south visitcentraloncley
17 Wed 01-Mar-2006central servicecentraloncley
16 Tue 28-Feb-2006central boom anglescentraloncley

265: HotFilm, Site central, Sun 30-Apr-2006 17:33:22 PDT, Inverter failure ends hot-film run at 1000 LT 30 Apr
The power inverter lent by Oncley to power Ripple in the hot-film/stn0 electronics box was unable to operate due to low battery power of unknown origin. A brief effort to ascertain the genesis of the low power was made approximately 30 minutes prior to the end of T-REX (00UTC 1 May). We soon decided that the benefit was limited given the end of the project (we had originally intended to run until the Central tower was shut down).

The hot-films were gathered by Poulos at ~ 1900 UTC 30 Apr.

Thanks for all the hard work and the gathering of some 300 hours of data.


262: power, Site central, Sun 30-Apr-2006 10:26:37 PDT, H2 level
Hydrogen level is at 900psi at 1345 local time on 4/30/06.
258: LOG, Site central, Sat 29-Apr-2006 15:35:01 PDT, ISFF media day
On late 28 Apr 2006 Poulos and Golubieski took the Haines-Stiles videographer to the Central site. In addition to larger questions regarding the overall success of T-REX an extensive set of high resolution (digital camera quality) video was taken of the Central tower and the OTIHS 5m tower nearby. Detailed pictures of the hot-film anemometers, motors and operations were taken. Kevin said that these were 'cool'. An amusing experience to say the least.
256: LOG, Site central, Fri 28-Apr-2006 17:22:32 PDT, Central Guy Tensions
Tensions on 30M tower

NW guys (inner - outer): 400,340,310,300
SE guys (inner - outer): 340,260,250,300
SW guys (inner - outer): 400,320,275,275

Taken approx. 1630 local time.

255: krypton, Site central, Fri 28-Apr-2006 17:19:52 PDT, Krypton Cleaning
30M Krypton clean 1615 LT.

243: wireless-network, Site central, Tue 25-Apr-2006 18:22:29 PDT, Wireless fiddle tests
Poulos and Agee executed tests where the Central and Hot-film etherants were switched between HotWire and NIDS SSIDs. We also manually altered the elevation and azimuth of the high gain antenna. 

In all cases while occasional improvements to dB were found (as high as 18dB), the system dB always varied significantly and to low levels (down to 6-8dB). No preferential orientation could be systematically established.

It was noted that REAL and West continued to maintain 20dB+ levels.

The DLR antenna has been returned and we could try it on Central since it consistently returned 20dB+ as well, but given the late state of the experiment we will likely be satisfied with reboots of the Central etherant 12.40 until the end.
242: wireless-network, Site central, Tue 25-Apr-2006 15:23:21 PDT, Wireless intermittent outages, return to the original configuration 12.40
Overnight the performance of the recently implemented eantJ 12.42 at the Central tower was flaky. Although initially a steady performing antenna with between 14-16dB, this system gradually reduced in effectiveness and was found at ~10dB this morning.

The plan then was to move this antenna to the high-gain antenna, which, with little hot-film traffic would be relatively free for Central. In addition, the high-gain antenna could increase the SNR of the new Central antenna by as much as 6dB ideally. With John's help Greg implemented new SSID that was incorrect and shutdown the wireless link to Central. We used 'hotfilm' instead of the actual name 'HotWire' and rebooted unsuccessfully.

As a result the central tower was visited and the ToughBook brought up into the ethernet board, which showed that 12.42 could not be pinged. An attempt was made to try a ping from the Central NIDS to no avail. The system was hardbooted in an attempt to recover functionality, unsuccessfully. As a result some data was entirely lost (roughly 45 minutes). 

Data began coming in again after switching the cable to the original antenna, 12.40, eantH, which had been left up on the tower disconnected when eantJ (12.42) was put up, just in case.

We found that 12.40 came right up with activity in all three lights, whereas reconnecting 12.42 there was 2 lights flickering and the wireless activity light not on. This indicates a wireless configuration problem, I suspect.

12.42 was detached and brought back to base. We do not plan to work on reconfiguring it here now since 12.40 is operating with 12-14dB currently.

We have not attempted to use HotWire as SSID with 12.40 yet.

240: HotFilm, Site central, Mon 24-Apr-2006 19:28:23 PDT, Hotfilm run
24Apr06, 17:40PDT, Central
Began hotfilm test..

End. 25 Apr 2006 ~230pm local.

Dual backup of data completed 26 Apr 2006

239: krypton, Site central, Mon 24-Apr-2006 19:27:03 PDT, Cleaned 5m Krypton
24Apr06, 17:30PDT, Central

Cleaned 5m krypton on short tower
238: rain, Site central, Mon 24-Apr-2006 14:56:19 PDT, Rain Check, Download Data
24Apr06, ~13:45PDT, Central

Yesterday we had 
a) good rain 
b) lots of missing data from central (see other logs)
so darn, we couldn't see the rain quantities!

Today we checked the rain gauge.  It had roughly a residual of ~1/3 bucket,
and we did a tip test (offline from NIDS)......OK, working.

So we halted the viper and swapped in the 4GB usb stick to bring in and
merge the local storage from yesterday....more later...
Yup, the events are there.

237: soil, Site central, Mon 24-Apr-2006 14:49:36 PDT, Tsoil Coefficients, etc.
24Apr06, 13:40PDT, Central

Greg has noticed that the Tsoil readings have appeared quite low during the
entire experiment.   Steve Oncley reviewed the readings and agreed.  Discussion
with Steve Semmer suggested the possibility that the coefficients in the logger
could be incorrect, so we checked:

Sensor Installed per label on cable= STP97001, same sensor as spreadsheet
indicates it should be.
Coefficients were exactly as what the 'excel-spreadsheet' said they should be:
Tbl4-Location	Value		Spreadsheet
30		-244.0		-244
31		+235.89		235.89
32		+7.6499		7.65
The spreadsheet file is in/called:
-rw-r--r--  1 aster users 31744 Apr  6 16:00 radiationinfo.xls

Re: wiring
A subsequent check and the wires appeared OK.  I did notice that the E3
excitation seemed a bit odd to me: 4-each 5mS pulses spaced at 5mS apart
and at the correct 2.1VDC.

236: soil, Site central, Mon 24-Apr-2006 14:39:55 PDT, Trime Measurements, Gravimetric
24Apr06, 13:30, Central

~14:15PDT soil sample taken
Definitely moist from rain yesterday, so good opportunity.  Seems quite well
formed and clay-like except for the 'loose alkaline top.'

0-3cm	125.0-45.1 = 79.9gm
	e1	25.8	20.0
	e2	15.8	12.4
	e3	18.5	14.4

3-6cm	169.8-73.1 = 96.7gm
	d1	20.2	14.9
	d2	21.3	15.7
	d3	18.5	13.6

14:00PDT Trime Greg took these readings:
Sample	%Vol	TDR
1	39.8	24
2	42.3	13
3	39.8	0
4	39.6	24
5	42.1	12
6	39.8	0
7	39.6	24
8	42.1	12
9	39.1	0
10	38.2	22
11	38.7	22
12	39.3	22

All 4 sets of 3 readings had good penetration at 45-deg angle in the vicinity
of the station soil sensors.
#1-3, #4-6, #7-9 were in separate locations.  Because of the funny 0 TDR
values as the measurements proceded, the last set #10-12 were taken.
Ground definitely has the moisture in it.

232: power, Site central, Sun 23-Apr-2006 19:43:02 PDT, Hydrogen replaced at Central
The H2 at Central was replaced with a full (2200PSI) container. This will last until the last day of the experiment.
228: LOG, Site central, Sat 22-Apr-2006 16:04:43 PDT, Boom angles, central
The boom angles at Central were shot on Saturday morning and Sunday morning (2 and 5 m).

Sonic  Solar   Solar   Targeted     Boom     Corrtn     Boom angle
 Hgt    Time    Azim    Value       Angle    to angl  after correction
------ ------ ------- ------------ -------   -------- ----------------
  1 Sat 1120  134.25d  134d 35.7m  88d 42m    -20.7m   88d 21.3m
  2 Sun 1005  112.11d  112d 27.5m  87d 47.5m  -20.9m   87d 26.6m
  3 Sat 1120  134.25d  134d 35.7m  88d 42m    -20.7m   88d 21.3m
  5 Sun 1025  116.92d  117d 03m    85d 18.5m   -7.8m   85d 10.7m
 10 Sat 1045  122.69d  122d 20.5m  92d 28m    +20.9m   92d 48.9m
 15 Sat 0955  110.21d  110d 44m    92d 59m    -31.4m   92d 27.6m
 20 Sat 1100  127.29d  127d 18m    93d 52.5m   -0.6m   93d 51.9m
 25 Sat 0955  110.21d  110d 44m    92d 59m    -31.4m   92d 27.6m
 30 Sat 1100  127.29d  127d 18m    93d 52.5m   -0.6m   93d 51.9m

225: wireless-network, Site central, Sat 22-Apr-2006 13:58:33 PDT, Outages continue; New Eant at Central
22Apr06, 8:30-9:30 PDT

The central etherant was down a couple times yesterday and was restarted twice
as late as midnight.  Then it was down again this morning at 5am.
Yesterday I pulled the eantj-sodar off the sodar to swap in at central.

When I climbed to central EantH, here are a couple observations while it
was still 'down.'
- power and wireless activity lights were present, looked good.
- the ethernet connection light was out.
	Note, Greg had wanted to go through swapping the pocketec proceduer
	before doing the antenna, and we had some troubles again
	with getting the swap-in to go.  However the ethernet should have
	been up, esp after a hard reboot of the viper.
	This could indicate a problem perhaps between antenna and
	viper instead of to tpop....maybe...
- the cable end going into the Eant, is improper because there is a gap
	between where the insulation was stripped back and the backshell
	of the connector.
	I taped this connector up speculating that it may be an outside
	possibility that our troubles began after the rain in late Mar,
	with perhaps some moisture getting in?....unlikely.
- 15vdc was OK.
Swapped in EantJ-sodar, installed just below the 'H'.   I pointed it slightly
differently, aiming a bit more down and a bit more to the north.....uncertain
whether this 'adjustment' might help or not.   It may move the pointing more to
the side lobe and perhaps away from interfering sources in town??
Left EantH installed in case it proves to be better than the new one.

Re: SNR seen from the Eants
From briefly monitoring the webpage for EantH yesterday we were saw ~13-15dbSNR,
which is marginal (ie 'yellow').  This afternoon while watching the EantJ,
we've been noticing anywhere from 9-17db reported from the antenna with
nominal around 12.

Re: Using EantJ-sodar instead of EantG-school at central.
I put EantG-school on the sodar yesterday afternoon instead of using it for
central.  The reason is because last summer while doing our range tests, 'G'
had been flakey and had an 'X' penciled on it.  During preparations for TREX
it did work OK in the staging area so decided to ship it hoping it would be
used for 'the other guys' or as a last resort.  Instead it found its way onto
the school, a critical link.   It's not out of the question that some of our
problems with 'School-network' were due to 'G' being flakey??
Meanwhile since yesterday afternoon it seems to be happy at the sodar site
200yds from the base.

223: HotFilm, Site central, Fri 21-Apr-2006 14:03:20 PDT, GPS coordinates of the hot-film sonic anemometers
GPS coords under the 1m sonic
222: licor-co2, Site central, Fri 21-Apr-2006 14:01:15 PDT, Pictures around Licor
Poulos took three pictures in the vicinity of the 5m tower near Central where the Licor CO2 is sited ~ 1330pm 21 Apr.
221: LOG, Site central, Fri 21-Apr-2006 11:39:07 PDT, Central pocketec download
Central data from 4/11-4/21 was downloaded to exthd7 on aster on 4/21.

The pocketec from the hot-film system is now at central as local storage.

219: HotFilm, Site central, Thu 20-Apr-2006 19:25:28 PDT, hotfilm running

17:20PDT began new hotfilm run with 'fresh' pocketec disk.
218: krypton, Site central, Thu 20-Apr-2006 19:24:31 PDT, 5m Krypton cleaned
20Apr06,, 17:10PDT

Cleaned with methyl alcohol
213: soil, Site central, Wed 19-Apr-2006 19:41:43 PDT, Trime Measurement, Gravimetric
19Apr06, Central

12:20PDT soil sample taken
Some roots in the sample.

0-3cm	126.0-79.3=46.7gm
	e1	11.6	10.6
	e2	15.2	14.0
	e3	14.9	13.6

3-6cm	151.8-73.1=88.7gm
	d1	20.7	16.1
	d2	16.1	12.5
	d3	20.6	15.9

11:40PDT Trime:
Sample	%Vol	TDR
1	42	16
2	41.7	16
3	44.4	13
4	45.1	13
5	37.2	15
6	36.7	15

All 3 sets of 2 readings had good penetration at 45-deg angle in the vicinity
of the station soil sensors.
Measurements 5,6 were taken closer to a bunch of grass on the north side.

212: radiation, Site central, Wed 19-Apr-2006 13:33:31 PDT, Rad.cleaned
19Apr06, 12:25-33

Cleaned all domes.  The Rnet and some of the long, short wave radiometers did show a bit of dirt on the tissue.  The Rnet had visible dust on its domes but not heavy.

204: HotFilm, Site central, Tue 18-Apr-2006 09:54:16 PDT, hotfilm system clock ahead
ntp on daisy was mis-configured, and as a result its
clock was  591.58 seconds ahead.

This was corrected at 09:19 this morning, and the
system rebooted, so data recorded after this time are OK.

I believe it has been mis-configured since it was
put on the 192.168.15 network early in the project.

201: HotFilm, Site central, Sun 16-Apr-2006 19:32:39 PDT, Started hotfilm
Started hotfilm sampling at 19:30 PDT, Apr 16.

This run lasted until 20Apr06 ~12:30 when it was taken down for awhile to
allow the full pocketec to be off-loaded at the base
197: HotFilm, Site central, Sat 15-Apr-2006 17:32:25 PDT, hotfilm power
ripple (hotfilm Dell notebook) died this afternoon

It's internal battery was discharged. The inverter that provides
power was giving a little beep every 5 seconds or so and the green
led on the inverter would brighten and then dim.

Looks like the voltage at the 5m tower is not high enough to drive
the inverter.  The two batteries that were at the tower were moved
to the 30m tower this morning.  I moved one back over to the 5m
tower and then the inverter was happy.  However, then the main
power breaker on the enclosure would flip. ripple's internal
battery was discharged, so it was pulling it's highest load.

Installed a power 'Y' so that the power to the inverter is coming
straight from the battery, rather then via the aux plug on the
enclosure.   Seems to work. It is running right now.

196: data-system, Site central, Sat 15-Apr-2006 12:59:55 PDT, Short local storage outage
15Apr06, 11:15-12:06PDT

When I reconfigured power at central and booted the DSM back up, the
pocketec drive failed start up properly.
At 12:05 I cycled power on the pocketec and it started up again.

195: power, Site central, Sat 15-Apr-2006 12:47:08 PDT, Reconfigured power to hotfilm tower.
15Apr06, ~11:00-11:20

Took down ripple and the hotfilm system, plus central tower to
reconfigure the power distribution so that both the main tower and the
hotfilm components are being operated through the 'Iload' circuit.
This way we can monitor our 'full' consumption.

At ~13.4vdc the current peaked at roughly 10.8A when the hotfilm
system was booted up.  After that it has settled down to roughly 6.5A.
So we can expect perhaps 7-8A when the voltage drops.  I checked the
pam_reference document and confirmed the trace is rated 10A and
installed a big fuse so it won't blow.  For overnight at least I
unplugged the beacon.  It seems that we don't really need it since the
site is clost to the much taller cell tower at this location, plus FAA
requirements are 200'.

The ripple/inverter are now running off 'Aux1' on the hotfilm power
distribution panel after the fuse was replaced.  Some care will be
needed when booting up the hotfilm box.  Probably best to:
1) unplug Daisy inside box.
2) unplug Ripple/inverter from Aux1
3) turn on box and observe motors initialize
4) boot Ripple
5) boot Daisy

194: radiation, Site central, Sat 15-Apr-2006 12:44:09 PDT, 10m Rad cleaned, photos
15Apr06, 11:00PDT
Cleaned 10m Rad.  No visible evidence of dust.
A full set of 360, etc photos were taken from the 10m perspective.

193: TRH, Site central, Sat 15-Apr-2006 12:38:43 PDT, 5m TRH filter check
15Apr06, 10:30PDT
The 5m TRH was off for a few minutes while inspecting its filter.
Good news, it appeared very clean although there was a bit of this fine dust on the support piece.
If this site's TRH is clean then almost certainly the rest are also.
191: HotFilm, Site central, Fri 14-Apr-2006 15:29:32 PDT, hotfilm data system test
Another hot-film data system test was started at ~ 1420pm (without hot-films themselves) during the latest Central visit.
190: power, Site central, Fri 14-Apr-2006 15:16:23 PDT, Central 2m sonic, power usage
John/Greg visited Central at about 1230 and replaced the 2m sonic cable and confirmed working status at ~ 1425pm.
During this visit we ran a series of checks on power consumption and found the following:
The power adapter for Ripple in the Daisy box draws 0.3A
Ripple, with the screen active and the power adapter, draws ~ 1.5A but somewhat more during certain activities.
The total system draws about 1.5A pulsing to 2+ A with disk writes (includes Ripple, Daisy, hot-film motors, etherant, CTAs and the Air-adapter).

All of Central was found to draw ~6A during the day without the beacon. The 35W fuel cell continued to provide 2A. The Lind Auto-air Adapter was found to be making beeps, which we'll investigate.
186: HotFilm, Site central, Thu 13-Apr-2006 18:05:34 PDT, hotfilm data system test
Hotfilm data system is up - but no probes are installed.

The prometheus is running a modified version 5.91 of the DSCUD 
driver and library.

The old version was 5.9.  I doubt 5.91 has any bug fixes - most
likely it just includes support for a new board from Diamond.

I've made a slight modification to the 5.91 driver - declaring
as volatile two variables that are shared between the 
interrupt routine and the user read method.  The A2D stayed
up all last night and today but we didn't always have sonics
connected, and never had motors connected.

14Apr06 6:45PDT  Ripple halted
Gordon had to use a 4gb usbstick instead of the pocketec for this test.  Since the ingest was down due to full disk, I halted ripple to save power since it's cloudy and the batteries are down to 12.15
185: sonic, Site central, Thu 13-Apr-2006 18:00:30 PDT, 2m Sonic down
Cable for 2m sonic at central failed when the hotfilm system
was being re-installed. The 2m sonic is now only connected to
182: power, Site central, Thu 13-Apr-2006 12:36:21 PDT, Update on 35W Central Fuel Cell
Central 13Apr06, 10:30 PDT
Last night we said...
(after putting 35W on main batt bank) We'll observe overnight and expect some charging for central out of this unit.  If it works, we may move the fuel cell and batteries adjacent to the rest of the stuff and use the long cable when needed to run the hotfilm system.

Overnight it did provide 1A charging despite the long cable and blocking diode.
Today we moved it to the main battery bank, and removed the blocking diode going into the Icharge2 circuit.

Currently the config. is the solar is charging through the prostar on Icharge1,
and 35W f.c. charging through the '8A' charger in it's box, on Icharge2

The fuel cell then began charging at ~2A, and the solar at ~13.5A
swapped tank in:
2500psi, regulator set at 15psi to f.c.
181: sonic, Site central, Thu 13-Apr-2006 12:34:38 PDT, Guy Wires added to 5m tower
13Apr06, ~10:30PDT

Installed 2 'cross' guy wires to 5m tower to try to help stablizing the oscillations in the wind.

NOTE: this may have effected the sonic levels as the cables were tightened.
180: krypton, Site central, Thu 13-Apr-2006 12:19:17 PDT, 30m krypton cleaned
13Apr06, ~10:00PDT

- 30m Krypton cleaned.  Voltage went from ~.9 to ~1.1 vdc

178: power, Site central, Wed 12-Apr-2006 11:37:10 PDT, Update on Central Fuel Cell / Battery Status
Central 12Apr06, 9:00-14:30 PDT
Last night we said...
"For overnight we'll run the fuel cell straight onto the battery and hope it comes to life."

It never really did get going and at most provided .5A charging.  This morning the tank has about 2400psi.  We fiddled with the controls, pressures, etc. some more but decided this '60W' fuel cell requires more effort than is advisable in the field plus since it isn't doing anything, and given the note several have made about it's 2 'questionable' internal cells, we removed it from central and brought it back to the base.

35W fuel cell put on central, ~9:50PDT
Since Greg reports no hotfilm runs are imminent, we reconfigured to put the 'happy' 35W f.c. through the central charge monitor 'Icharge2' circuit in place of the bad unit.  This is going through a long (~100') cable, and the blocking diode and then straight to the battery at the end of the chain.  We still have the 2 batteries tied directly on at the 35W f.c.   Given the voltage drop of the cable and the diode, plus the solar charging in full sun we are seeing:
.85A charging from 35-f.c.
13.4 VDC at batteries on f.c.
12.7 VDC at battery on Icharge2
1100 psi remaining in tank.

We'll observe overnight and expect some charging for central out of this unit.  If it works, we may move the fuel cell and batteries adjacent to the rest of the stuff and use the long cable when needed to run the hotfilm system.
177: radiation, Site central, Wed 12-Apr-2006 11:33:03 PDT, 'cleaned' 10m K&Z
12Apr06, ~10:15
attempted to clean the radiometer with methanol using Steve Semmer's sponge mop.
The wind was substantial so I only got the downward looking dome swiped.  There was no evidence of any dirt either by visual inspection or by looking at the 'swabber.'  Consequently I did not maneauver around the logger, etc. to swipe the upper looking dome.   The suspicion is there is just too much wind for this very fine dust to stick, and not enough moisture.   We have not seen any rain.
176: krypton, Site central, Wed 12-Apr-2006 11:31:02 PDT, 5m krypton cleaned
12Apr06, ~9:40 PDT
cleaned 5m krypton on short tower with methanol.
Voltage changed from ~.2 to ~.5, not spectacular.

Note: very windy and this short tower swaps a great deal.

174: power, Site central, Tue 11-Apr-2006 15:41:20 PDT, Update on Central Fuel Cell / Battery Status
Notes for 2-days regarding the fuel cell:

Central 10Apr06, 14:00-14:30 PDT
Installed New charge controller with SOAD onto /ttyS13
Swapped in new H2 bottle: 2500psi, regulated at 20psi
Current from f.c. no more than .15A

Note: we did observe a very brief time where the fuel cell 'claimed' to have provided about 4Amps charging, but only for a few minutes.  After that it stopped.

It was very very nice to finally have the SOAD reporting Iload, vbatt, Icharge1 (pv) and Icharge2 (cell).....thanks Gordon!

11Apr06, 13-14PDT
H2 bottle down to 900psi.  No charging overnight. VOC=~16
Is it simply purging the bottles??

In theory that going through prostar charge controller may be a problem, we installed a segregated set of wires through Icharge2 directly to the end of the battery chain, avoiding the controller entirely.  We also swapped in a new H2 bottle.  After restarting we observed on both the SOAD and directly on a meter ~.45A charging to the battery.

NOTE: the cell-cell voltages showed most at ~.75v, but one was at -.15 and another was fluctuating from ~.2-.75.
Possible that these 'bad' cells are draining the current exactly like a bad battery cell would do.  Need to measure their impedances.

NOTE: We did observe a slight, but noticable increase of roughly .15+ A out when we openned the lid of the electronics box enclosure.

For overnight we'll run the fuel cell straight onto the battery and hope it comes to life.

171: HotFilm, Site central, Mon 10-Apr-2006 19:59:30 PDT, Hotfilm run
Begin 10Apr06,19:45 PDT Started hotfilm.
End   11Apr06 12:15 PDT

The fuel cell running ripple is looking good: 1200psi remaining, and the 2 batteries are at 13.6 on the f.c. before we started.  At the end, the batteries were still 13.6 and the f.c. 1100psi.
The main system batteries were at 12.55, and thus ~11.55 going into daisy at the end of the long cable when we began.

168: power, Site central, Sun 09-Apr-2006 18:36:56 PDT, Update on Central Fuel Cell / Battery Status
The Batteries at Central have been gradually charging with the full sun and the hotfilm system off most of the time.

Fuel Cell System at Central
Note: this note begins from 7Apr when I did a check after dark.

19:40 PDT, after dark.
	Fuel Cell VOC     = 14.4	seems OK
	Fuel Cell Icharge = 0		bad. despite batteries at 12.5
	Metered open ground on Icharge2
	Moved Fuel Cell to 'solar-in'
	Fuel Cell Icharge = .12A	yuck. thus the non-consumption of H2
					the last several days.
	0) Central has been running only on solar.  The solar system here is
		just enough based on manual battery monitoring the last few
		days.  The added load of daisy/ethant/hotfilm puts it on the
	1) The Fuel Cell is basically only keeping it's fans running at this
		point.  The filter masks are in place but noticably dirty.
		It's possible if I'd left the f.c. connected to the solar-in
		port longer that it may have 'started-in' but that's T.B.D.
	2) The charge monitor board isn't wired properly to allow the F.C.
		to work even if it were running ok.

	The next day, Apr8th, we removed the charge monitor that has an SOAD
	from the fuel cell box and brought it to the base to get running.  After
	making a cable for the SOAD - NIDS we confirmed the serial 
	interface/command and Gordon began work on 'polling software' for NIDS.
	This will allow us to use the SOAD instead of fudging in a logger for
	the power monitor, plus gives another important NIDS/EVE software
	capability.  By Sunday evening Gordon had this code working on the
	test box.
	We 'fixed' the charge-monitor board so the 2nd charging supply (f.c.)
	can work and added the blocking diode.  Intend to install tomorrow.

165: radiation, Site central, Sun 09-Apr-2006 14:33:50 PDT, Radiometers 'Cleaned'
11:30PDT 4/9/06 Cleaned radiometer domes.
These were all in good shape and did not appear to have any 'dirt' on them.
Note: Did not clean 10m rad

161: HotFilm, Site central, Sun 09-Apr-2006 09:05:57 PDT, HotFilm Run Sa. 4/8/06 - Su. 4/9/06
4/8/06 ~17:00PDT Begin
4/9/06 ~11:30PDT End of run

Fuel Cell began at 1400psi for ripple supply.
Fuel Cell end   at 1250psi, Vbatt for 2 running ripple=13.5 on F.C.
Central power system for Daisy, etc. was looking good also after 3 sunny days of solar charging.

160: soil, Site central, Fri 07-Apr-2006 18:21:23 PDT, gravimetric soil moisture
Took soil sample at central at ~15:50 PDT, April 7

I intentionally took this sample near where the probes enter the ground and a sage bush.  There were dry stems and roots in this sample, esp in the 0-3 level.  These were for the most part removed for the 'shake-n-bake'

0-3 cm
Total weight = 70.3 - 42.9 gm = 27.4 gm
	wet	dry
b1	6.2	5.4
b2	8.5	7.4
b3	11.5	10.1
Note: I think this 0-3cm is bogus because: 1) procedural error on my part, 2) very dry soil and high winds blew it while transferring.   I.e. don't believe the bulk because of 'suspicious volume'.... this one might be for practise only.

3-6 cm
Total weight = 177.1 - 73.0 = 104.1 gm
	wet	dry
e1	26.0	20.8
e2	21.8	17.4
e3	20.0	15.9
Note: This one was better and should be OK.  It did have some roots and a small worm that needed rescuing before cooking the batch.

159: power, Site central, Fri 07-Apr-2006 17:15:48 PDT, Central Battery Status
Battery voltage checks at Central today:
We have had full sunshine yesterday and today so the charging system should be running full tilt.
~ 11:30 PDT  VDC = 12.60 at 'end' battery
~ 15:30 PDT  VDC = 12.66 at 'end' battery
~ 18:00 PDT  VDC = 12.95 at battery adjacent charger
	Sun going down in sky.
	Icharge-Solar = 7.5A going in
	Icharge-F.C.  = 0.0A
	500psi on F.C. in am.
The end battery is running the daisy/motors so has had a bit higher load the last few days.

~ 19:40 PDT  VDC = 12.45  after Shutdown Daisy/Ripple

158: HotFilm, Site central, Fri 07-Apr-2006 17:09:54 PDT, Imitation Hotfilm Test Running
Winds picking up today.
Begin ~22:30Z

No hotfilms installed and CTA's not running but otherwise daisy/motors/ripple are up so the Gordon can install some new scripts for improved data recovery over the periodic reboot method.

Note: Fuel cell was still running OK.  By that time the batteries were at 12.7 on the f.c. charging.   Swapped in 2 fully charged batteries.

157: HotFilm, Site central, Fri 07-Apr-2006 17:01:08 PDT, Hotfilm Shutdown
Winds picking up today.
Stop Test: Apr 7th, 097: 18:35Z
Started: Apr 5th, 

Note: Fuel Cell was barely running when I arrived.  The regulator showed 0-psi on both gauges.  The battery voltages were at 12.2 at that point.  When I checked the valve it was barely cracked open.  When I openned it all the way the pressure went to 1600psi and etc.  At that point I had shutdown everything.

155: rain, Site central, Thu 06-Apr-2006 10:55:28 PDT, no rain data at central
The rain gauge at central has been connected to the wrong channel on the 
campbell data logger.  This was fixed this morning.

154: wireless-network, Site central, Thu 06-Apr-2006 10:45:41 PDT, wireless network from central down
The wireless network from central was down from about 9:25 - 10:35 PDT.
Steve and John restarted by cycling power on the etherant at central.
152: HotFilm, Site central, Wed 05-Apr-2006 18:34:19 PDT, Hotfilm running
The hotfilm is up and running at ~18:00.

151: wireless-network, Site central, Wed 05-Apr-2006 18:02:14 PDT, central communications lost
Wireless communication to central lost around 9 am PDT, restored around
1800 PDT.
150: radiation, Site central, Tue 04-Apr-2006 16:09:33 PDT, ZigBee now power on/off at Central
The two single pygreometer logger boxes, #1 and #2, now have their radios
being turned on then off by the logger. They are setup to be on for 1 second
and off for 4 seconds. New batteries were insatlled in each unit. The
code change can be found on the Panasonic under the PC208W directory.
The file name is TREXzig.

147: wireless-network, Site central, Tue 04-Apr-2006 09:50:39 PDT, lost rf connection to central
Lost wireless connection to central around 0600 PST this morning.  Briefly lost
connection to west while Steve trying to fix this problem and dsm server and 
netcdf process stopped.  Fixed connection to west and restarted dsm server and
netcdf process.  Steve recycled power on the etherant at central to restart 
around 10 am PDT.

145: wireless-network, Site central, Tue 04-Apr-2006 07:57:43 PDT, Link to central down
The link to central has gone down. This occurred at about 6:00 am this morning.
The link to the hotfilm is still good so I do not think it is the TPOP;
however, I will recycle power on TPOP to see if that fixes it.
Power recycle did not help. Time for visit to central.

142: LOG, Site central, Mon 03-Apr-2006 15:39:48 PDT, retensioned guy wires
Tightened the N and SE, highest guy wires by two turns.  Could not tighten
the SW guy wire because the turnbuckle is fully compressed.

After tensioning and with a strong south wind:

N = 290 lbs
SW = 340 
SE = 320
141: HotFilm, Site central, Mon 03-Apr-2006 10:56:05 PDT, hot film procedures
Hot Film:
	The hot film system consists of 4 major components: (1) power source, 
(2) sensors (hot film and sonics), (3) data collection (Prometheus and Ripple), 
and (4) the communication system. The system can run without the communication 
link going however you will not be able to monitor on a real time basis.

     This assumes that power is off on all systems and all sensor cables are 
plugged in.

1: Power system
     The power system is designed to be running at all times. The power system 
consists of 2 parts; the batteries and the fuel cell. The fuel cell keeps the 
batteries charged.  Verify that the fuel cell system is running. If it is not 
operating then check the hydrogen tank. The tank pressure should be at least 
300 psi. If it is below 300 psi then it should be changed.  If you do change 
the tank remember what the setting of the regulator is  between 10 and 20 psi 
on the second stage. After the tank is attached then you can start the fuel 
cell. There is an on/off switch on the fuel cell located below the hydrogen 
line. Turn the switch on and the fuel cell should start. It will make a noise 
for about 15 seconds then start running.  If it stops you may need to replace 
the 9 volt battery located in the fuel cell. This requires removing the fuel 
cell from the enclosure. The battery is located on the side of the cell.

2:  Startup Ripple
     Ripple is located inside the electronics box attached to the top of the 
lid. Open the box and plug the power cable into one of the power ports on the 
12 volt battery splitter cable. Attach the USB drive if it is not connected. 
This is necessary in order for the aster processes to get going. Undo the 
bungee cord so you can open the DELL. BE CAREFUL UNHOOKING THE BUNGEE CORD. 
Power up the DELL and login when requested. 
Do a check_aster command to verify that the aster process is running. 
Also do a ping aster to verify that the wireless communication link is up.

3: motor controllers
     The motor control systems should be plugged into serial ports 9 (1m), 
10 (2m), and 11 (3m). With the lid open verify that the internal power cord to 
the Prometheus stack is unplugged. Now turn the main power switch on. This 
should start the 3 motor systems. The zero position sequence starts with the 
motor shaft moving CW to about 135 degrees. It stops for a second then goes 
CCW to the zero position. Zero is parallel to the sonic boom direction. You 
want to make sure all three probe support arms follow the sequence above. If 
something does not go right then turn off the main power switch immediately.
NOTE: At this time NO probes should be mounted.

4: Tri-axial probes
	The tri-axial probes can be installed. Note the reference number on 
each probe box. The probe order is: #1 – 1m, #2 – 2m, #4 – 3m. When 
inserting the probes note the notch cut out at the back end of the probe. 
This should be aligned with a corresponding piece on the probe support. 
Be extra careful handling the probes!  Put the yellow protection cap, 
on the end of the support arms into the probe box.
5: Prometheus
	The Prometheus can now be plugged in. You can run minicom on ripple 
if you want to watch it boot up. After a minute or two you should see a cockpit 
come up on ripple. Verify everything is running as expected. Also verify data 
is being written to the USB drive. This can be done by either watching for a 
green light flash or do a DF command on ripple and verify that it is increasing.
At this time the hot film data on the analog channels should be around zero 

6: CTA power
	Now plug in the CTA power cable. It is plugged into on of the aux ports
on electronics box.  Move it over to the other aux port. You should be getting 
A/D on the cockpit.

	Shutdown is pretty much in the reverse order as startup

1: CTA power
	Unplug the CTA power cord and move it back to the aux port.

2: Prometheus
	Via minicom or ssh to daisy, do a halt command to stop the Prometheus.

3: Tri-axial probes
	Remove the hot film probes. Do them one at a time to ensure they get 
put back in their correct box. 
Remember to put back on the yellow protection cap.

4: Ripple
	Ripple can now be shutdown. Once it is shutdown remove the USB drive so
that the data can be downloaded back at the base. 
Remember to unplug the power cable and put it back inside the box.

Other items of interest:
Sonics: the sonics are routed to 2 locations: Prometheus and central. 
Only the transmit signal is routed to the Prometheus. If you ever have to 
disconnect the sonic cables from the Prometheus box make sure you note how 
they are connected.
Sonic 1m – port 5
Sonic 2m – port 8
Sonic 3m – port 7

Ripple: Ripple may be running. Tests have shown that the fuel cell is apply to 
keep ripple up all the time if its internal battery is fully charged. However, 
if the USB drive is not attached the aster processes may not be running!

136: HotFilm, Site central, Sun 02-Apr-2006 08:04:11 PDT, Hot film status
The hot film system was up and running around 16:00 yesterday. The fuel cell
system was able to keep ripple going since it was powered up for Gordon on
The deadband was set to 20 degrees.

134: sonic, Site central, Sat 01-Apr-2006 15:11:28 PST, 25 m sonic at central
Restarted 25 m sonic with rserial -h /dev/ttyS10 central.
Then & T Ae D & (?)
132: krypton, Site central, Sat 01-Apr-2006 13:48:14 PST, maintenance at central
Steve and Tom at central tower.  Steve climbed 30 m tower 11:40-12:20 PST to
investigate communication problem and clean kh2o at 30 m.  Then cleaned kh2o
and Licor 7500 at 5 m; off 5 m tower at 12:34 PST.

KH2O s/n = 1394 @ 30 m
           1390 @ 5 m

Finally, Steve cleaned 10 m radiometers with mop.
Also added yokes for steps and safety rope on 10 m radiometer tower.
127: HotFilm, Site central, Thu 30-Mar-2006 11:37:01 PST, HotFilm shutoff
The hotfilm was shutdown at about 10:45am local. The sonics were moved over
to central nids.
The 35 watt fuel cell kept things going and the batteries still are up
around 12;5 volts.

124: HotFilm, Site central, Wed 29-Mar-2006 16:55:35 PST, Hot Film status
The hot film was offically back on the air at 16:35.
During the process the 1m, 2m, and 3m sonics were removed from nids central
and move back to daisy. An option to eliminate this cable changing is to
build special Y cables to route the sonic xmt line into daisy. Maybe in the

123: radiation, Site central, Wed 29-Mar-2006 16:53:13 PST, Logger #1 at central
The logger # 1 at central is back up (~16:15). The code changes to turn on
and off the ZigBee did not work. The transceiver comes on but the receive
light does not. The logger was put back into tis old mode.

121: radiation, Site central, Wed 29-Mar-2006 10:18:24 PST, Shutdown pyg sensor at central, logger ID 2
I have removed logger ID 2 from central to make some modes to the software
so that the radio will go to sleep between data transmissions. This should
help the power issue at the two single pygreometer sites, The logger was
removed around 9:15 am local,

119: HotFilm, Site central, Tue 28-Mar-2006 14:58:41 PST, Hot film status
The hot film system was left up overnight to collect sonics data.
Unfortunately the batteries could not keep things up and the system went down.
Earlier this morning I changed the batteries and brought the system back up.
I re-routed power for daisy to come from the nids power source. The fuel cell
system will now provide power for ripple only.

The sytem was taking down once we got the sonics interfaced to nids (refer to
messages 117 and 118). We will see if the fuel cell keeps the batteries
charged at hot film.

118: sonic, Site central, Tue 28-Mar-2006 14:55:57 PST, Hot film sonics on central nids
The 3 hotfilm sonics are now going through nids central. This was done
by extending the sonic cables to each the nids box. Whenever the hot film
is active these sonics need to be reconnected to Daisy.
117: sonic, Site central, Tue 28-Mar-2006 13:19:05 PST, hot-film sonics added to NIDS
In order to collect the 1,2,3 m sonic data when the hot-films are not running,
we are cabling them to the viper data system.

I have saved ads3.xml as ds3.xml.old and modified ads3.xml with the
following additional code:


115: HotFilm, Site central, Mon 27-Mar-2006 15:42:17 PST, HF site back up collecting sonic data at 15:30
The hotfilm site was brought back on the air collecting sonci data only.
This was done vias channel_config. There are now wo channel_configs in
OPS. "channel_config.sonci" is used for only sonic data while
"channel_config.hf" is the hotfilms added. You have to copy either of
these files to "channel_config" before starting things up.

The battery voltage levels were down around 11.75 volts. The hydrogen
tank was at ~1800 psi. This will be a good test to see if the fuel cell
can recharge the batteries wihtout the hot film stuff going. I also
disconnected power to the CTA moduels.

113: radiation, Site central, Mon 27-Mar-2006 11:42:15 PST, Battery changed at logger #1 at central
The battery was changed at radiaiotn logger #1 at central.
This took place around 10:45 this morning.

112: HotFilm, Site central, Mon 27-Mar-2006 11:26:34 PST, Hot film status
The hot film ran through out the night until ~10:30 this morning. The data
are bieng downloaded to exthd7 in the base. Once the pocketec drive is
cleared off it will be takne back out to central where the system will be
turned on to collect sonic data only.
NOTE: the battery level was down to 12 volts. The fuel cell was running
fine so this down time will be good for the batteries.

110: HotFilm, Site central, Sun 26-Mar-2006 14:10:07 PST, Deploying hot film probes
Deployed all three hot film probes this afternoon for a night run.

Data started about 1500 pst.

106: HotFilm, Site central, Sat 25-Mar-2006 17:04:09 PST, hotfilm update
Mods have been completed on the motor controller circuit. All 3 units are
now deployed at central. The plan is to do a run tonight with breakable
wooden sticks in place of the probes. If the motor controllers go beyond
the limit switches the wooden probes should break.
If things run goo over night then we are back in business to collect hotfilm

103: HotFilm, Site central, Fri 24-Mar-2006 17:30:25 PST, hotfilm status
Work was spent on the motor controller systems. SInce one of the optic units
was not acting correctly I decided to re-wire it to a different format. There
is now a pull down resistor on the emitter side of the detector. The
result was very positive. The signal levels pull all the way to 0 volts as
wanted. I will need to re-write the code because the trigger logic is now
reversed; ie, it triggers on a falling edge verses a rising edge. I should
get this done tonight so that we can deploy them tomorrow.

101: wireless-network, Site central, Fri 24-Mar-2006 10:51:16 PST, data link to central fixed
The data link to central went down around mid-day yesterday; when Steve was
climbing tower??

Went to central this morning & recyled power on both the etherant and the viper.
Link working again.
99: HotFilm, Site central, Fri 24-Mar-2006 07:51:04 PST, Hotfilm work from yesterday
The hotfilm system was shut down yesterday. I brought the motor controller
systems back to the base. We are having a problem with the coupler between
the motor shaft and the vertical probe support shaft. Also problems with the
coupler and the optics units.

97: radiation, Site central, Thu 23-Mar-2006 13:47:19 PST, maintenance at central site
March 23: Steve and Tom
Service visit to central site
Guywire tensions:
1	380	450	420
2	325	310	350
3	310	370	300
4	320	300	305

Steve cleaned KH2O and LI-7500 with methanol
Steve on 30 m tower 11:35 - 12:05
Steve of 5 m tower  12:05 - 12:10

Then cleaned radiometers with methanol spray, wiped off radiometers
with kleenex.

94: HotFilm, Site central, Thu 23-Mar-2006 07:45:27 PST, hotfilm
Based on the archive files on ripple, diasy shut down around 3:45 this morning.
I assume that the reset program "ndaqrestart" used in the crontab may have a
bug. I will talk to Gordon.

I plan to go out there in teh next 1/2 hour to shut the system down and bring
back the USB drive.

91: HotFilm, Site central, Wed 22-Mar-2006 17:44:30 PST, htofilm - probe position problem
All shafts on probe position system have loosened up. I believe the problem
is the tolerences between the optics and the shaft coupler along with the pin
used to trigger the optic. I have seen two problems:
(1) The shaft coupler piece has had contact with an optic detector causing
the detector to move out of position.
(2) The roll pin has made contact with the optic detector as it passes through.
The contact point has been at the bottom of the optic detector.

Possible solution:
(1) reduce the diameter of the shaft coupler towiden the space between the
coupler and the optics.
(2) change the roll pin to a flat surface piece. This would provide more
clearence between the top and bottom of the optic piece.
90: HotFilm, Site central, Wed 22-Mar-2006 17:31:54 PST, Installed hot film sensors
Installed hot film probes and began running at ~1712 pst.

Had problems with orientation shafts slipping, so turned off orientation motors
and oriented probes ~ west.

1 m probe had problem with y bnc connector immediately after probe.  Steve
jiggered with it and got it connected for now. I believ this is the connector
that John had found to be bad.

The batteries were changed. The old olds were done to ~12 volts.The fuel
cell should be able to recharge the batteries during the daytime when
the system is off. 

Probe 1 at 1 m
Probe 2 at 2 m
Probe 4 at 3 m

87: HotFilm, Site central, Tue 21-Mar-2006 16:55:56 PST, hotfilm
   I believe the problem with ripple has been found. It may have been in the USB
port connection. The PCMCIA USB card is a USB2 connection while the USB
connection on the back of ripple is a USB1. I changed to the ripple USB at
about 11:30 this morning and it has been running fine every since.

  Prometheus is still having problems. It only wants to stay up for a few
minutes. The plan now is to install a forced reboot on Prometheus every xx
minutes where xx will be determined based on how it is running.

   The fuelcell is working like it was meant to be out there. It has used about
100 PSI over the last 24 hours.

86: HotFilm, Site central, Tue 21-Mar-2006 10:11:27 PST, hotfilm status
During the night ripple shut down so no data was collected. The 35 watt
fuel cell kept runnin gok and the batteries are charged. I went out this
morning to restart ripple. I also remounted the 1m motor system. I had to
remove the 2m motor system becasue the motor shaft coupler screws were loose.

83: HotFilm, Site central, Mon 20-Mar-2006 16:47:31 PST, Htofilm status
During the day the hotfilm was left running at central. The A/D did not stay
up all the time. Back at the base I got the 35 watt fuel cell running. It 
was taken out to central to serve as the main power source for the hotfim
stuff. Tis included ripple which is mounted inside the Prometheus box.
In addition Gordon made some changes so that the USB harddrive is now
interfaced to ripple instead of the Prometheus. The system was back on the
air at about 16:40 local time.

79: HotFilm, Site central, Mon 20-Mar-2006 07:30:27 PST, hotfilm
The system stayed up during the night. Based on the local storage file sizes
the A/D stayed up for about 7 hours. Ripple stayed up on an external battery
that is not on the charging grid. I will go out there later this morning to
check on the ripple battery and restart the A/D system.

78: power, Site central, Sun 19-Mar-2006 18:35:04 PST, fuel cell
The 60 watt fuel cell was at 14.68 volts and 800 psi at 6:15pm.

77: HotFilm, Site central, Sun 19-Mar-2006 18:29:42 PST, hotfilm status
The hotfilm was moved out to central this afternoon. There were some initial
crashes on the A/D. After some playing around it seemed to stay up. The 3m
sonic was not responding correctly. Who knows if this may have been having
some effect. It was disconnected. Ripple is running off an external battery
via the car charging unit that Steve Oncley had. There are no hotfilm probes
deployed. Also the motor controller at the 1m height was acting up. It was
not seeing the zero position switch. It appears to be a similar problem we
were having back in Boulder. What is real frustrating is that unit was in the
base during the tests over the last few days and it worked without any

74: power, Site central, Sat 18-Mar-2006 14:15:48 PST, central fuel cell
At noon the fuel cell was reading 15.3 volts and the pressure was at 1050.

73: LOG, Site central, Sat 18-Mar-2006 14:13:47 PST, Central visit
Tom and I went to central to do a few tasks. This included getting a soil sample, relocating the solar panel a logger 1 and checking the fuel cell. This took place around noon.

71: power, Site central, Sat 18-Mar-2006 08:37:34 PST, Fuel Cell at central
I did a check on the fuel cell this morning.
At 8:20 voltage = 14.80 gas = 1100

69: HotFilm, Site central, Fri 17-Mar-2006 17:49:31 PST, HotFilm
Due to the discovery of the daisychain on the sonics, any data looked at since this time should be done with great care. In fact any sonic data representing 2m and 3m should be ignored.
68: HotFilm, Site central, Fri 17-Mar-2006 15:15:53 PST, Removed hot film data system and 1 m sonic
Steve S and Tom removed the hot film data system and 1 m sonic from central
site this afternoon to work on it in lab.

Steve discovered that the 2 and 3 m sonics were never cabled within the data 
system.  The 3-sonic replicator cable was still transmitting the 1 m sonic 
data to all three channels.It was also discovered that serial channel 201 was
not responding to a sonic input. So a change was made in channels.txt on daisy
(/home/isff). Channel 201 was moved from /dev/ttyS5 to /dev/ttyS7

Later that evening:
  Steve set the Hotfilm system up in the trailer. It has been running for
~1hour with one sonic and 9 analog at 2000 hz. It is now runing with 3
sonics and 9 analog. It will be left up over night to see what happens.

67: power, Site central, Fri 17-Mar-2006 10:59:42 PST, 60 watt fuel cell
I checked the output of the 60 watt fuel cell around 10:k00 am. It was at about 13 volts. So just for fun I tiurned down the fan power and the cell jumped up to about 15.5 volts. I plan on going out there at sundown to monitor it.

66: radiation, Site central, Fri 17-Mar-2006 10:55:58 PST, single pyg units
Spent some time out at central changing the batteries and re-orienting the solar panels at the two single pyg sites. This took place between 9:00 and 10:30 am.

62: LOG, Site central, Thu 16-Mar-2006 14:56:48 PST, Central visit
- Replaced empty H2 cylinder with full.
- Cylinder we're running on is at 1360psi -- about twice the usage compared to
the last 2 nights.

59: HotFilm, Site central, Thu 16-Mar-2006 09:31:00 PST, prometheus config change
As yet another test, I've just changed the config to enable sampling of only
1 triple hotfilm probe at 1000 and all 3 sonics.  The usb disk didn't come up.
I had tried to set the sample rate to 500 on all hotfilms, but the prometheus
went into a continuous ndaq restarting cycle -- perhaps 500 sps is not a valid

I had to start archiving manually, which is now being done to ripple's disk.
I figured that we might as well be acquiring data from 3 sonics, since they
don't significantly add to the data rate.

P.S. Even this test was a failure!  It ran for only 20 minutes and died.  
Data rates in the file were ok (no dropouts -- frequency as expected).

57: power, Site central, Wed 15-Mar-2006 14:45:57 PST, Service visit to Central tower
Tom and Steve visited the Central tower after the West tower.

Check guywire tensions:

SE upper guys: upper guy = 300 lbs	325 lbs
               lower guy = 320        	315 
N  upper guys: upper guy = 280		300
               lower guy = 370		370
SW upper guys: upper guy = 300		325
               lower guy = 325		350

Second reading above is after increasing the tension on the uppermost guy wire
by one full turn.

H2 pressure = 2000 psi

Steve on tower cleaning 30 m KH2O: 12:05 - 12:20 PST
Steve on 5 m tower cleaning 5 m KH2O: 12:25 - 12:30 PST

Apparently cleaning the 30 m KH2O made little change in the output, but cleaning
the 5 m KH2O decreased the output from 0.9 V to 0.7 V !!

54: HotFilm, Site central, Wed 15-Mar-2006 08:49:48 PST, hotfilm analogs died
cockpit shows that the A/D process on the hotfilm system died sometime between
John and I leaving at about 7 last night and us arriving about 8:45 this 
morning.  I'll leave it going, since the sonics (only the lowest 2 are enabled)
are still running.

53: power, Site central, Tue 14-Mar-2006 18:46:37 PST, cylinder readings
When we connected the fuel cell to the full cylinder yesterday morning, it was
reading 2500psi.  This morning, the reading was 2260psi -- a normal change,
for once!

50: radiation, Site central, Tue 14-Mar-2006 11:57:26 PST, Tcase working at central
After several conversations, Steve S/John/I found a programming error in the
central 4-comp logger.  The P7 instruction needed single-ended channel 3, not
2.  Made the change manually (*7h*1...*0) and cycled power to check that it
was saved.  Now Tcase ratios are 1.35, 1.34, and 1.34 for a1, a2, and 4-comp
"out"s, respectively!  The 10m values are different, but perhaps this is 
real and/or a calibration issue.

47: radiation, Site central, Mon 13-Mar-2006 12:21:25 PST, wiring error
Found that the wire logger L2 to Amp pin 6 had been inserted in pin 8 by 
mistake.  Fixed this about 11:30.  Didn't seem to help the Tcase data :(

45: data-system, Site central, Sun 12-Mar-2006 16:51:17 PST, central died last night
Central died last night presumably due to power (we added the load of a 
second etherant plus the prometheus yesterday evening).  Came up by itself in
the morning -- probably about 1 hour lost.  Also rebooted during the day 
when new batteries (which hopefully will help the power problem) were installed
at about 1pm(?).  It appears that the local usb kept up with these cycles.

As usual, lost 15m sonic data until manually recycled both times.

43: HotFilm, Site central, Sun 12-Mar-2006 16:37:43 PST, hotfilm stuff installed
Spent much of today at central (John/Greg/ sometimes Steve&Brad) installing
the hotfilm stuff.  Left it with middle (nominally 2m) wire installed, but
motors not working at about 3:10.

Also checked tensions -- somewhat low.  We need to retension sometime.

39: radiation, Site central, Sat 11-Mar-2006 11:12:54 PST, visit
Was at central ~10:15 to:
- check fuel cell: 13.45V/1260psi
- clean all rads (including 10m for the first time)
- TRIME: 38%/11%/28% (from west to east) -- basically unusable.
- check logger cabling.  Only difference between a2 and 4-comp on K&Z is white
wire goes to "AG" on a2 rather than "G" on 4-comp.  TP01 wiring looks the
same as I remember at west.

38: power, Site central, Sat 11-Mar-2006 09:07:24 PST, added fuel cell
Added the 60W to central (with rather jury-rigged cabling) at about 18:00 last
night.  Didn't seem to do much, i.e. didn't seem like the controller was 
putting a load on it, even though it was dusk.  Cylinder pressure ~1250.  Fuel 
cell voltage output 15.55V.

We'll go out now and check up on it.

36: radiation, Site central, Fri 10-Mar-2006 11:01:39 PST, Radiometer Power Issues
Steve and I spent this morning trying to figure out why two of the radiation stands at central and the rad stand at west were not working.  A simple fix to get them running while we troubleshoot the problem was to switch out the batteries with ones we had in the trailer.  There was a decent amount of dust on the solar panels at central, but we are unsure how much of a factor it is.  
28: power, Site central, Mon 06-Mar-2006 12:17:52 PST, central power
Central stayed up overnight.  Today is overcast and some drizzle.
Measured all 5 batteries (disconnected from each other, though 1 was kept tied
to the station)
1	12.32
2	12.41
3	12.43
4	12.42
5	12.45 (the one I added yesterday)
All seem reasonable, though Kurt thinks they should be stronger and is planning
to get new batteries to us sometime.

18: LOG, Site central, Wed 01-Mar-2006 14:00:30 PST, south visit
South visit ~11:30-12:40

- cleaned rads
- cleaned kryptons
- TRIME soil moisture 3.9/6.8/5.2%

17: LOG, Site central, Wed 01-Mar-2006 13:55:37 PST, central service
Visited central ~1000-1100
- 10 tips on rain gauge at 1002
- cleaned rads (except 10m)
- TRIME readings: gave "salinity too high", "no TDR signal", eventually got
readings of 20/23/40%.  (Use 40%??.)
- cleaned kryptons (5m is a hard reach)
- tried to level 5m csat better, but boom always droops (I suspect no pin in
- diagnosed csat 5m problem as cable with a short.  Will replace on the way
to lunch
- measured aux1 voltage 13.24V

16: sonic, Site central, Tue 28-Feb-2006 19:51:07 PST, central boom angles
Shot boom angles today, though the theodolite base wouldn't turn much -- had
to do a VERY coarse adjustment rotating the entire theodolite head.

Fortunately, much of the tower looked the same:

30m	93 deg 13'
25m	92 deg 35'
20m	92 deg 35'
15m	92 deg 35'
10m	92 deg 35'
5m	85 deg 21' [separate tower]